find the point estimate of the population parameter μ, and the standard deviation of the given set of data below;


Percentage of children who watch TV before bedtime

70 67 58 60 69 62 70 62

69 59 77 59 52 79 59 59

80 42 60 59 68 40 68 68

56 66 60 40 57 57 70 71

72 54 52 67 62 59 71 72

81 49 45 78 78 69 68 69



1
Expert's answer
2021-06-06T16:37:57-0400

point estimate of the population parameter μ Mean:

"\mu=\frac{\sum x_i}{N}"


"\mu=\frac{70+67+58+60+...+68+69
}{48}=\frac{3039}{48}"

"=63.3125"


Standard deviation:

"\sigma=\sqrt{\frac{\sum(x_i-\mu)^2}{N}}"


"\\sigma=\\sqrt{\\frac{\t\n(70 - 63.3125)2 + ... + (69 - 63.3125)2}{48}}=10.048126380077"